# Natural polynomials for Kerr quasi-normal modes

### v2

@@ Abstract
We present a polynomial basis that exactly tridiagonalizes Teukolsky’s radial equation for quasi-normal modes. These polynomials naturally emerge from the radial problem, and they are “canonical” in that they possess key features of classical polynomials. Our canonical polynomials may be constructed using various methods, the simplest of which is the Gram-Schmidt process. In contrast with other polynomial bases, our polynomials allow for Teukolsky’s radial equation to be represented as a simple matrix eigenvalue equation that has well-behaved asymptotics and is free of non-physical solutions. We expect that our polynomials will be useful for better understanding the Kerr quasinormal modes’ properties, particularly their prospective spatial completeness and orthogonality. We show that our polynomials are closely related to the confluent Heun and Pollaczek-Jacobi type polynomials. Consequently, our construction of polynomials may be used to tridiagonalize other instances of the confluent Heun equation. We apply our polynomials to a series of simple examples, including: (1) the high accuracy numerical computation of radial eigenvalues, (2) the evaluation and validation of quasinormal mode solutions to Teukolsky’s radial equation, and (3) the use of Schwarzschild radial functions to represent those of Kerr. Along the way, a potentially new concept, “confluent Heun polynomial/non-polynomial duality”, is encountered and applied to show that some quasinormal mode separation constants are well approximated by confluent Heun polynomial eigenvalues. We briefly discuss the implications of our results on various topics, including the prospective spatial completeness of Kerr quasinormal modes.
@@ I Introduction
The results of single black hole (BH) perturbation theory have proved widely useful to gravitational wave astronomy Vishveshwara (1970); Abbott et al. (2016a, b, 2021a, 2021b, 2019); Hughes et al. (2019); Khan et al. (2016); Husa et al. (2016); Pompili et al. (2023). Most famously, the aftermath of binary black hole (BBH) coalescence is expected to be a perturbed Kerr BH that rings down, radiating away its perturbation, predominantly in the form of damped oscillatory modes Press (1971); Press and Teukolsky (1973); Teukolsky (1973); Teukolsky and Press (1974); Leaver (1985); Berti et al. (2006a); London et al. (2014); Hamilton et al. (2023). The related ringdown radiation is of practical relevance to many aspects of gravitational wave astronomy, from the construction of signal models, to the ongoing search for new physics London et al. (2014); Hamilton et al. (2023); Hughes et al. (2019); Khan et al. (2016); Husa et al. (2016); Pompili et al. (2023); Carullo et al. (2018); Baibhav et al. (2023); Carullo et al. (2019); Cotesta et al. (2022); Mishra et al. (2010); Bhagwat et al. (2020); Abbott et al. (2016b). Concurrently, ringdown radiation has inspired varied activity in gravitational wave theory, from efforts to understand BH nonlinearity, to emerging developments in time-dependent perturbation theory Lagos and Hui (2022); Green et al. (2023); Cannizzaro et al. (2023); Sberna et al. (2022); Redondo-Yuste et al. (2023). These efforts, among many others, aim to prepare gravitational wave science for a future in which high-precision ringdown measurements may point the way to a more complete understanding of gravity Arun et al. (2022); Maselli et al. (2017); Berti et al. (2016); Barausse et al. (2015); Klein et al. (2016); Bailes et al. (2021); Vallisneri et al. (2015); Karnesis et al. (2022); Arun et al. (2022). There is, therefore, good reason to think that a deeper practical understanding of linearly perturbed Kerr BHs is of use.
@@
Here, the overtone index is shown to originate from an order parameter inherent to the confluent Heun polynomials Ronveaux and Arscott (1995); DLMF ; Fiziev (2009a); Hortacsu (2012); Dariescu et al. (2021); Magnus et al. (2021). Concurrently, it is observed perhaps for the first time that the confluent Heun polynomials exhibit a kind of “polynomial/non-polynomial duality”, where finite sets of confluent Heun polynomials are necessarily paired with an infinite set of non-polynomial functions.
@@
–  It was pointed out in Ref. London (2023a) (hereafter Paper I) that there exists a scalar product for which Teukolsky’s radial equation is self-adjoint, and that this scalar product enables the direct application of Sturm-Liouville theory Pinchover and Rubinstein (2005); Abramowitz and Stegun (1964); Arfken et al. (2013); Courant and Hilbert. (1954). The scalar product’s weight function coincides with the Green’s function approach to QNMs pioneered by Leaver, and the deformed Jacobi weight known for Pollaczek–Jacobi polynomialsLondon (2023a); Leaver (1986); Chen and Dai (2010). One use of the scalar product is to represent Teukolsky’s radial equation as a simple linear matrix equation Axler (2015).
While many studies have developed matrix representations of Teukolsky’s radial problem, the lack of an appropriate scalar product has caused complications, such as the need to numerically integrate the equations, or appearance of non-physical modes Chung et al. (2023a); Blázquez-Salcedo et al. (2023); Chung et al. (2023b); Ghojogh et al. (2023); Zhu et al. (2023). In contrast, the scalar product presented in Paper I was hypothesized to result in matrix representations that are explicitly symmetric. Further, it was postulated that there may exist a basis of classical-like polynomials that enable the representation of Teukolsky’s radial equation as a symmetric and tridiagonal matrix. Such matrices are known to have number of favorable properties; e.g., their eigensolutions contain no non-physical modes, and they are well developed methods to study their asymptotic structure Teschl (2000); Kuijlaars et al. (2003).
Here, it is shown there do indeed exist unique classical-like polynomials that exactly tridiagonalize Teukolsky’s radial equation. Therefore, the resulting method of solution to the QNM’s radial problem is spectral in the simplest possible way: widely available linear algebra packages may be used to explicitly diagonalize the problem, simultaneously yielding eigenvalues and eigenfunctions. The polynomials in question will be referred to as “canonical confluent Heun polynomials”; numerical examples are provided.
@@
Here, the QNMs’ radial functions are shown to be generally non-polynomial in nature, but potentially well approximated by a confluent Heun polynomial of order given by $n+1$, or in at least one case ($n=8$), of an order much greater. A simple procedure for quantifying the approximate polynomial nature of a QNM’s radial function is presented.
@@
–  There has been a significant amount of research into the question of QNM completeness222 As most commonly defined, a function sequence is complete (i.e. is a basis) if any piecewise continuous function on the same domain can be exactly expressed as a linear combination of sequence elements. It is a standard result in Sturm-Liouville theory and functional analysis that self-adjoint differential operators have complete eigenfunctions.. Some references appear at tension regarding whether the radial field equation is self-adjoint, and therefore whether or not QNMs radial functions may be complete. This apparent tension arises from the fact it is insufficient to a label a differential equation as non-self-adjoint, without additional statements about the coordinate domain, boundary conditions, and related bilinear form (e.g. inner or scalar product). For example, Refs. Jaramillo et al. (2021); Leung et al. (1994); Nollert (1999) describe analogs of the radial field equation as non-self-adjoint; however, akin to Refs. Wald (2008); Beyer (1995); Green et al. (2023), a key result of Paper I was the construction of a radial scalar product such that the equation is self-adjoint.
@@ II Scope and organization
–  The focus of this article will be on linear order curvature perturbations of isolated Kerr BHs, as framed by Teukolsky, Press and Leaver Teukolsky (1973); Press and Teukolsky (1973); Teukolsky and Press (1974); Leaver (1985, 1986). While this article discusses only gravitational perturbations (i.e. those with spin-weight $s=\pm 2$), all calculations present make no explicit assumption about spin-weight, and in most cases have been verified to explicitly work with other physically relevant spin weights, namely $|s|\leq 2$. The authors do not aspire to the language of formal mathematical rigor; in place of theorems and lemmas, arguments are presented as needed, and the most technical of these are left to references. Numerical results will use QNM frequencies due to their physical relevance; however, this article’s results use a framework that applies to any single frequency. The concurrent treatment of all QNM frequencies is left to future work. The unit convention, $G=c=1$, will be used throughout.
@@
–  This article is organized as follows. In Sec. III, a schematic background of single BH perturbations is provided, culminating in the definition of the QNM’s radial eigenvalue problem, and a brief review of Paper I. In Sec. IV, the QNM’s radial eigenvalue problem is treated with confluent Heun polynomials. This section concludes with a discussion of the confluent Heun polynomials pros and cons. In Sec. V, this article’s central result, namely the canonical confluent Heun polynomials, are introduced. In Sec. VI, the canonical confluent Heun polynomials are shown to tridiagonalize Teukolsky’s radial equation. In Sec. VII, numerical results are presented to demonstrate the canonical confluent Heun polynomials’ end-to-end use. In Sec. VIII, the results of the preceding sections are discussed in the context of past and future work.
@@ III Preliminaries
–  A gravitational wave observatory will detect a linear combination of the two strain polarizations present in Einstein’s General Relativity (GR), $h_{+}$ and $h_{\times}$ Blanchet (2014). We may parameterize each using Boyer-Lindquist coordinates, $\{t,r,\theta,\phi\}$ where, asymptotically far from a radiating source, $t$ may be defined to measure time elapsed for an observer, $r$ the luminosity distance, and $(\theta,\phi)$ the usual spherical polar angles Boyer and Lindquist (2004). Since all $\theta$ dependence is in the form of $u=\cos(\theta)$, $u$ will henceforth be used Leaver (1985).
@@
[…]
In Eq. (1b), $\psi_{4}$ is the fifth scalar of the Weyl tensor in the Newman-Penrose formalism Newman and Penrose (1962). It is a field of spin-weight $s=-2$, and encodes all relevant information about gravitational radiation Newman and Penrose (1962, 1966); Ruiz et al. (2008); Merlin et al. (2016). For a single isolated BH, the rescaled Weyl scalar
@@
satisfies Teukolsky’s equation Teukolsky (1973); Press and Teukolsky (1973); Teukolsky and Press (1974). Schematically, if we denote the second order linear partial differential operator within Teukolsky’s equation as $\mathcal{L}_{tru\phi}$, then for isolated black holes,
[…]
@@
In Eq. (4), $\tilde{\omega}\;=\;\omega\,-\,i/\tau$ is complex valued QNM or “ringdown” frequency, and $\tau$ must be positive for time domain stability of the BH Press and Teukolsky (1973); Leaver (1985); Whiting (1989). Consequently, in the time domain, each QNM is a damped sinusoid with central frequency $\omega$, and damping time $\tau$. Teukolsky’s equation constrains the functional form of each QNM, and the details of the perturber constrain its overall excitation amplitude Leaver (1986); Berti and Cardoso (2006).
@@
Henceforth, Eq. (5a) and Eq. (5b) will be referred to as Teukolsky’s angular and radial equation, respectively. They are eigenvalue equations, where $-A$ and $+A$ are the respective complex valued [^3] The complex nature of the eigenvalues is a consequence of the fact that Eqs. (5a) and (5b) have complex valued terms. It will be seen in Eq. (14) that this coincides with a non-conjugate symmetric scalar product. eigenvalues London (2023a, b); Fiziev (2009b); Leaver (1986). Therefore, $S(u)$ and $R(r)$ are the respective eigenfunctions. In Eq. (III), $\mathcal{L}_{u}$ and $\mathcal{L}_{r}$ are second order linear differential operators that will be referred to as Teukolsky’s angular and radial operators, respectively. Each implicitly depends on $s$, $m$ and $\tilde{\omega}$, as well as the BH mass $M$, and spin $a=|\vec{J}|/M$.
@@
Figure 1: Distributions of radial (black dots) and angular eigenvalues (open circles) for $s=-2$, BH dimensionless spin $a/M=0.7$, and the $(\ell,m,n)=(4,1,3)$ QNM with $M\tilde{\omega}_{413}=0.8545-0.6305i$. Since both eigenvalue distributions have been computed with a QNM frequency, they have one coincident value corresponding to $(\ell^{\prime},n^{\prime})=(\ell,n)=(4,3)$. Radial egenvalues shown have been computed using the results of the current work, and verified using Leaver’s method Leaver (1985).
$\mathcal{L}^{\prime}{{}_{u}}\,S_{{\ell^{\prime}}}(u)\;$ $=\;-A_{{\ell^{\prime}}}\,S_{{\ell^{\prime}}}(u)\;,$ (6a)
$\mathcal{L}^{\prime}{{}_{r}}\,R_{{n^{\prime}}}(r)\;$ $=\;+A_{{n^{\prime}}}\,R_{{n^{\prime}}}(r)\;.$ (6b)
Equations (6a) and (6b) hold for any frequency parameter $\tilde{\omega}^{\prime}$ that is independent of $\ell^{\prime}$ and $n^{\prime}$. Since $\mathcal{L}{{}_{u}}$ and $\mathcal{L}{{}_{r}}$ depend on the frequency parameter, they are also primed in Eq. (III). It is only for the QNM frequencies that $A_{\ell^{\prime}}=A_{n^{\prime}}=A$, and Eq. (III) is recovered. Since $\ell^{\prime}$ and $n^{\prime}$ are independent, for any one QNM frequency, there will be only one value of $\ell^{\prime}$ and one value of $n^{\prime}$ such that $A_{\ell^{\prime}}=A_{n^{\prime}}$. These special labels are those for which $\ell^{\prime}=\ell$ and $n^{\prime}=n$. For these reasons, the QNM frequencies and their separation constants are often labeled with spatial mode numbers, i.e. $\tilde{\omega}_{{\ell mn}}{}$ and $A_{{\ell mn}}$ respectively Berti et al. (2006a); London et al. (2014); Berti et al. (2016); Magaña Zertuche et al. (2021). As a result, the QNMs are a non-trivial infinite sequence, wherein each element satisfies Eq. (III) with a nominally distinct frequency parameter, $\tilde{\omega}^{\prime}=\tilde{\omega}_{{\ell mn}}$ London (2023b).
An example of the angular and radial eigenvalue distributions is shown in Figure 1. There, the typical qualitative features of each distribution are shown: the angular eigenvalues, $A_{\ell^{\prime}}$, will typically be spaced horizontally along the real axis, and the radial eigenvalues, $A_{n^{\prime}}$, will vary most strongly in their imaginary part. Since the case shown has been constructed using a QNM frequency, there will be a single coincident eigenvalue, which is the one of physical interest (i.e. the one compatible with the construction of Eq. III).
@@
–  The angular equation, Eq. (6a), is well-known to be of the Sturm-Liouville type (i.e. $\mathcal{L}_{u}$ is self-adjoint) London (2023b, a); Courant and Hilbert. (1954); Lax (2002). Consequently, its eigenfunctions, $S_{\ell^{\prime}}(u)$, are complete and orthogonal with respect to a scalar product, ${\langle{a}\,|\,{b}\rangle}$, defined by the integral of $a(u)b(u)$ on $u\in[-1,1]$. The application of this angular scalar product is fundamental to the use of spherical and spheroidal harmonics in gravitational wave theory and data analysis Blanchet (2014); Khan et al. (2016); García-Quirós et al. (2020).
In contrast, the radial equation, Eq. (6b), is known to not explicitly be of the Sturm-Liouville type. This simply means that, in practice, the application of Sturm-Liouville theory must begin with the construction of a weighted scalar product, under which $\mathcal{L}_{r}$ is self-adjoint Courant and Hilbert. (1954); Pinchover and Rubinstein (2005); Kristensson (2010). Below, we provide a brief overview of the construction and evaluation of such a radial scalar product. We refer the reader to Paper I for a complete pedagogical development.
@@
Note that in Eq. (8), $\xi$ maps $r\in[0,\infty)$ to $\xi\in[0,1)$, thereby conpactifying the radial coordinate. In Eq. (9), $\mu(\xi)$ is comprised of factors related to the singular exponents of $\mathcal{L}_{r}$. Its explicit form is given in Eq. (12) of Paper I. In Eq. (III), $\text{C}_{0}$ through $\text{C}_{4}$ are complex valued constants determined by the BH mass $M$, the BH spin $a$, the mode numbers $\{\tilde{\omega},m,l,n\}$, and the field’s spin weight $s$:
@@
Together, Eqs. (8-III) allow the radial eigenvalue relation (Eq. 7) to be rewritten as
@@
[…]
@@
By combining the definition of $\mathcal{L}_{\xi}$ (Eq. III) with the requirement of self-adjointness (Eq. 15) and then applying integration by parts, it may be shown that
@@
In Eq. (19), $\xi^{j+k}\right>$ are called monomial moments Chihara (2011). Using the definition of the weight function, it has been observed (e.g. Cho et al. (2009); Leaver (1986); Abramowitz and Stegun (1964)) that each monomial moment $\xi^{p}\right>$ can be evaluated using the Gamma function, $\Gamma(x)$, and the Tricomi confluent hypergeometric function, $U(x,y,z)$,
$\xi^{p}\right>\;=\;e^{\text{B}_{2}}\,\Gamma($ $\text{B}_{0}+p+1)$
×$\times\;U(1+\text{B}_{0}+p,-\text{B}_{1},-\text{B}_{2})\;.$ (20)

### v3

@@ Abstract
We present a polynomial basis that exactly tridiagonalizes Teukolsky’s radial equation for quasi-normal modes. These polynomials naturally emerge from the radial problem, and they are “canonical” in that they possess key features of classical polynomials. Our canonical polynomials may be constructed using various methods, the simplest of which is the Gram-Schmidt process. In contrast with other polynomial bases, our polynomials allow for Teukolsky’s radial equation to be represented as a simple matrix eigenvalue equation. We expect that our polynomials will be useful for better understanding the Kerr quasinormal modes’ properties, particularly their prospective spatial completeness and orthogonality. We show that our polynomials are closely related to the confluent Heun and Pollaczek-Jacobi type polynomials. Consequently, our construction of polynomials may be used to tridiagonalize other instances of the confluent Heun equation. We apply our polynomials to a series of simple examples, including: (1) the high accuracy numerical computation of radial eigenvalues, (2) the evaluation and validation of quasinormal mode solutions to Teukolsky’s radial equation, and (3) the use of Schwarzschild radial functions to represent those of Kerr. Along the way, a potentially new concept, “polynomial/non-polynomial duality”, is encountered and applied to show that some quasinormal mode separation constants are well approximated by confluent Heun polynomial eigenvalues. We briefly discuss the implications of our results on various topics, including the prospective spatial completeness of Kerr quasinormal modes.
@@ I Introduction
The results of single black hole (BH) perturbation theory have proved widely useful to gravitational wave astronomy Vishveshwara (1970); Abbott et al. (2016a, b, 2021a, 2021b, 2019); Hughes et al. (2019); Khan et al. (2016); Husa et al. (2016); Pompili et al. (2023). Most famously, the aftermath of binary black hole (BBH) coalescence is expected to be a perturbed Kerr BH that rings down, radiating away its perturbation, predominantly in the form of damped oscillatory modes Press (1971); Press and Teukolsky (1973); Teukolsky (1973); Teukolsky and Press (1974); Leaver (1985); Berti et al. (2006a); London et al. (2014); Hamilton et al. (2023). The related ringdown radiation is of practical relevance to many aspects of gravitational wave astronomy, from the construction of signal models, to the ongoing search for new physics London et al. (2014); Hamilton et al. (2023); Hughes et al. (2019); Khan et al. (2016); Husa et al. (2016); Pompili et al. (2023); Carullo et al. (2018); Baibhav et al. (2023); Carullo et al. (2019); Cotesta et al. (2022); Mishra et al. (2010); Bhagwat et al. (2020); Abbott et al. (2016b). Concurrently, ringdown radiation has inspired varied activity in gravitational wave theory, from efforts to understand BH nonlinearity, to emerging developments in the time-dependent perturbation theory of single BHs Lagos and Hui (2022); Green et al. (2023); Cannizzaro et al. (2023); Sberna et al. (2022); Redondo-Yuste et al. (2023). These efforts, among many others, aim to prepare gravitational wave science for a future in which high-precision ringdown measurements may point the way to a more complete understanding of gravity Arun et al. (2022); Maselli et al. (2017); Berti et al. (2016); Barausse et al. (2015); Klein et al. (2016); Bailes et al. (2021); Vallisneri et al. (2015); Karnesis et al. (2022); Arun et al. (2022). There is, therefore, good reason to think that a deeper practical understanding of linearly perturbed Kerr BHs is of use.
@@
Here, the overtone index is demonstrated to be closely related to the order parameter inherent to the confluent Heun polynomials Ronveaux and Arscott (1995); DLMF ; Fiziev (2009a); Zarrinkamar and Hortaçsu (2018); Dariescu et al. (2021); Magnus et al. (2021). Concurrently, it is observed perhaps for the first time that the confluent Heun polynomials exhibit a kind of “polynomial/non-polynomial duality”, where finite sets of confluent Heun polynomials are necessarily paired with an infinite set of non-polynomial functions.
@@
–  It was pointed out in Ref. London (2023a) (hereafter Paper I) that there exists a scalar product for which Teukolsky’s radial equation, while non-Hermitian Rosas-Ortiz and Zelaya (2018); Mostafazadeh (2002); Andrianov et al. (2007); Berti et al. (2009b); Zhang et al. (2020); Brody (2013), is self-adjoint, and that this scalar product enables the direct application of Sturm-Liouville theory Pinchover and Rubinstein (2005); Abramowitz and Stegun (1964); Arfken et al. (2013); Courant and Hilbert (1989). The scalar product’s weight function coincides with the Green’s function approach to QNMs pioneered by Leaver [^2] The radial equation studied here is the same studied by Leaver Leaver (1985). Forthcoming numerical results are ultimately equivalent to those from Leaver’s series solution. However, truncations of Leaver’s series solution (the infinite dimensional problem) should not be confused with the canonical polynomials presented in Sec. V, which are used to frame a finite dimensional problem in Sec. VI., and the deformed Jacobi weight known for Pollaczek–Jacobi polynomialsLondon (2023a); Leaver (1986); Chen and Dai (2010). One use of the scalar product is to represent Teukolsky’s radial equation as a simple linear matrix equation Axler (2015).
While many studies have developed matrix representations of Teukolsky’s radial problem (e.g. Chung et al. (2023a); Blázquez-Salcedo et al. (2023); Chung et al. (2023b); Ghojogh et al. (2023); Zhu et al. (2023)), the scalar product presented in Paper I was is perhaps the first to result in matrix representations that are explicitly symmetric and tridiagonal, meaning that all physically relevant information is encoded in the diagonal and upper-diagonal matrix elements. Tridiagonal matrices are known to have number of favorable properties; e.g., they are well developed methods to study their asymptotic structure Teschl (2000); Kuijlaars et al. (2003); Chen and Dai (2010), and the classical properties of related polynomials enable efficient numerical methods Chen et al. (2019); Chen and Dai (2010); Chihara (2011). To arrive at a tridiagonal representation of Teukolsky’s radial equation, it was postulated that there may exist a basis of problem-specific, or “natural”, polynomials.
Here, it is numerically shown that there do indeed exist unique classical-like polynomials that exactly tridiagonalize Teukolsky’s radial equation. Therefore, the resulting method of solution to the QNM’s radial problem is spectral in the simplest possible way: widely available linear algebra packages may be used to explicitly diagonalize the problem, simultaneously yielding eigenvalues and eigenfunctions. The polynomials in question will be referred to as “canonical confluent Heun polynomials”; numerical examples are provided.
@@
Here, the QNMs’ radial functions are shown to be generally non-polynomial in nature, but potentially well approximated by a confluent Heun polynomial of order given by $n+1$. A simple procedure for quantifying the approximate polynomial nature of a QNM’s radial function is presented.
@@
–  There has been a significant amount of research into the question of QNM completeness [^3] A sequence of functions is complete (i.e. is a basis) if any piecewise continuous function on the same domain can be exactly expressed as a linear combination of sequence elements. It is a standard result in Sturm-Liouville theory and functional analysis that self-adjoint differential operators have complete eigenfunctions.. Some references appear at tension regarding whether the radial field equation is self-adjoint, and therefore whether or not QNMs radial functions may be complete. This apparent tension arises from the fact it is insufficient to a label a differential equation as non-self-adjoint, without additional statements about the coordinate domain, boundary conditions, and related bilinear form (e.g. inner or scalar product). For example, Refs. Jaramillo et al. (2021); Leung et al. (1994); Nollert (1999) describe analogs of the radial field equation as non-self-adjoint; however, akin to Refs. Wald (2008); Beyer (1995); Green et al. (2023), a key result of Paper I was the construction of a radial scalar product such that the equation is self-adjoint.
@@ II Scope and organization
–  The focus of this article will be on linear order curvature perturbations of isolated Kerr BHs, as framed by Teukolsky, Press and Leaver Teukolsky (1973); Press and Teukolsky (1973); Teukolsky and Press (1974); Leaver (1985, 1986). The related version of Einstein’s equations, namely Teukolsky’s equation, is well known to be non-Hermitian Berti et al. (2009b); Leung et al. (1994), and in that context, this work comments on select aspects of problem that are shared with Hermitian problems: symmetric matrix representations, orthogonality, and the potential completeness of eigenspaces London (2023b); London and Hughes (2022); London (2023a); Green et al. (2023). While this article discusses only gravitational perturbations (i.e. those with spin-weight $s=\pm 2$), all calculations present make no explicit assumption about spin-weight, and in most cases have been verified to work with other physically relevant spin weights, namely $|s|\leq 2$. The authors do not aspire to the language of formal mathematical rigor; in place of theorems and lemmas, arguments are presented as needed, and the most technical of these are left to references. Numerical results will use QNM frequencies due to their physical relevance; however, this article’s results use a framework that applies to any single frequency. The concurrent treatment of all QNM frequencies is left to future work. The unit convention, $G=c=1$, will be used throughout.
@@
–  This article is organized as follows. In Sec. III, a schematic background of single BH perturbations is provided, culminating in the definition of the QNM’s radial eigenvalue problem, and a brief review of Paper I. In Sec. IV, the QNM’s radial eigenvalue problem is treated with confluent Heun polynomials. This section concludes with a discussion of the confluent Heun polynomials pros and cons. In Sec. V, this article’s central result, namely the canonical confluent Heun polynomials, are introduced. In Sec. VI, the canonical confluent Heun polynomials are shown to tridiagonalize Teukolsky’s radial equation. In Sec. VII, numerical results are presented to demonstrate the canonical confluent Heun polynomials’ end-to-end use. There, we note that, due to the known asymptotic properties of the QNMs radial functions’ series solution, most truncated matrix representations of the radial problem will result in spectral pollution Nollert (1999); Petropoulou and Velázquez (2014); Boulton (2016); Lewin and Séré (2008); Davies and Plum (2003); Martínez Adame (2007). In Sec. VIII, the results of the preceding sections are discussed in the context of past and future work.
@@ III Preliminaries
–  A gravitational wave observatory will detect a linear combination of the two strain polarizations present in Einstein’s General Relativity (GR), $h_{+}$ and $h_{\times}$ Blanchet (2014). We may parameterize each using Boyer-Lindquist coordinates, $\{t,r,\theta,\phi\}$ where, asymptotically far from a radiating source, $t$ may be defined to measure time elapsed for an observer, $r$ the luminosity distance, and $(\theta,\phi)$ the usual spherical polar angles Boyer and Lindquist (2004). Since all $\theta$ dependence is in the form of $\cos(\theta), uu$ will henceforth be used Leaver (1985).
@@
$=\;\int_{-\infty}^{t}\int_{-\infty}^{t^{\prime}}\psi_{4}(t^{\prime\prime},r,u,\phi)\;dt^{\prime\prime}dt^{\prime}\;.$ (1b)
In Eq. (1b), $\psi_{4}$ is the fifth scalar of the Weyl tensor in the Newman-Penrose formalism Newman and Penrose (1962). It is a field of spin-weight s=2s=-2, and encodes all relevant information about gravitational radiation Newman and Penrose (1962, 1966); Ruiz et al. (2008); Merlin et al. (2016). For a single isolated BH, the rescaled Weyl scalar
@@
satisfies Teukolsky’s equation Teukolsky (1973); Press and Teukolsky (1973); Teukolsky and Press (1974). In Eq. (2), $a$ is the BH spin magnitude divided by the BH mass. Schematically, if we denote the second order linear partial differential operator within Teukolsky’s equation as $\mathcal{L}_{tru\phi}$, then for isolated black holes,
$\mathcal{L}_{tru\phi}\;\psi\;=\;0\;.$ (3)
@@
In Eq. (4), $\tilde{\omega}\;=\;\omega\,-\,i/\tau$ is a complex valued QNM or “ringdown” frequency, and $\tau$ must be positive for time domain stability of the BH Press and Teukolsky (1973); Leaver (1985); Whiting (1989). Consequently, in the time domain, each QNM is a damped sinusoid with central frequency $\omega$, and damping time $\tau$. Teukolsky’s equation constrains the functional form of each QNM, and the details of the perturber constrain its overall excitation amplitude Leaver (1986); Berti and Cardoso (2006).
@@
Henceforth, Eq. (5a) and Eq. (5b) will be referred to as Teukolsky’s angular and radial equation, respectively. They are eigenvalue equations, where A-A and +A+A are the respective complex valued [^4] The complex nature of the eigenvalues is a consequence of the fact that Eqs. (5a) and (5b) have complex valued terms. It will be seen in Eq. (14) that this coincides with a non-conjugate symmetric scalar product. eigenvalues London (2023a, b); Fiziev (2009b); Leaver (1986). Therefore, $S(u)$ and $R(r)$ are the respective eigenfunctions. In Eq. (5), $\mathcal{L}_{u}$ and $\mathcal{L}_{r}$ are second order linear differential operators that will be referred to as Teukolsky’s angular and radial operators, respectively. Each implicitly depends on $s$, $m$ and $\tilde{\omega}$, as well as the BH mass $M$, and spin $a=|\vec{J}|/M$.
@@
Figure 1: Distributions of radial (black dots) and angular eigenvalues (open circles) for s=2s=-2, BH dimensionless spin $a/M=0.7$, and the $(\ell,m,n)=(4,1,3)$ QNM with $M\tilde{\omega}_{413}=0.8545-0.6305i$. Since both eigenvalue distributions have been computed with a QNM frequency, they have one coincident value corresponding to $(\ell^{\prime},n^{\prime})=(\ell,n)=(4,3)$. Radial egenvalues shown have been computed using the results of the current work, and verified using Leaver’s method Leaver (1985).
$\mathcal{L}^{\prime}\hskip-2.5pt{{}_{u}}\,S_{{\ell^{\prime}}}(u)\;$ $=\;-A_{{\ell^{\prime}}}\,S_{{\ell^{\prime}}}(u)\;,$ (6a)
$\mathcal{L}^{\prime}\hskip-2.5pt{{}_{r}}\,R_{{n^{\prime}}}(r)\;$ $=\;+A_{{n^{\prime}}}\,R_{{n^{\prime}}}(r)\;.$ (6b)
Equations (6a) and (6b) hold for any frequency parameter $\tilde{\omega}^{\prime}$ that is independent of $\ell^{\prime}$ and $n^{\prime}$. Since $\mathcal{L}{{}_{u}}$ and $\mathcal{L}{{}_{r}}$ depend on the frequency parameter, they are also primed in Eq. (6). It is only for the QNM frequencies that $A_{\ell^{\prime}}=A_{n^{\prime}}=A$, and Eq. (5) is recovered. Since $\ell^{\prime}$ and $n^{\prime}$ are independent, for any one QNM frequency, there will be only one value of $\ell^{\prime}$ and one value of $n^{\prime}$ such that $A_{\ell^{\prime}}=A_{n^{\prime}}$. These special labels are those for which $\ell^{\prime}=\ell$ and $n^{\prime}=n$. For these reasons, the QNM frequencies and their separation constants are often labeled with spatial mode numbers, i.e. $\tilde{\omega}_{{\ell mn}}{}$ and $A_{{\ell mn}}$ respectively Berti et al. (2006a); London et al. (2014); Berti et al. (2016); Magaña Zertuche et al. (2021). As a result, the QNMs are a non-trivial infinite sequence, wherein each element satisfies Eq. (6) with a nominally distinct frequency parameter, $\tilde{\omega}^{\prime}=\tilde{\omega}_{{\ell mn}}$ London (2023b).
An example of the angular and radial eigenvalue distributions is shown in Figure 1. There, the typical qualitative features of each distribution are shown: the angular eigenvalues, $A_{\ell^{\prime}}$, will typically be spaced horizontally along the real axis, and the radial eigenvalues, $A_{n^{\prime}}$, will vary most strongly in their imaginary part. Since the case shown has been constructed using a QNM frequency, there will be a single coincident eigenvalue, which is the one of physical interest (i.e. the one compatible with the construction of Eq. 5).
@@
–  The angular equation, Eq. (6a), is well-known to be of the Sturm-Liouville type (i.e. $\mathcal{L}_{u}$ is self-adjoint) London (2023b, a); Courant and Hilbert (1989); Lax (2002). Consequently, its eigenfunctions, $S_{\ell^{\prime}}(u)$, are complete [^5] See Ref. Finster and Smoller (2016) for a discussion of completeness in the context of single complex valued oblateness parameters (i.e. $a\tilde{\omega}$), and see Ref. London (2023b) for related discussion of completeness in the QNM context, where oblateness parameters differ from mode to mode. and orthogonal with respect to a scalar product, $\langle{a}\,|\,{b}\rangle$}, defined by the integral of $a(u)b(u)$ on $u\in[-1,1]$. The application of this angular scalar product is fundamental to the use of spherical and spheroidal harmonics in gravitational wave theory and data analysis Blanchet (2014); Khan et al. (2016); Garc (i).
In contrast, the radial equation, Eq. (6b), is known to not explicitly be of the Sturm-Liouville type. This simply means that, in practice, the application of Sturm-Liouville theory must begin with the construction of a weighted scalar product, under which $\mathcal{L}_{r}$ is self-adjoint Courant and Hilbert (1989); Pinchover and Rubinstein (2005); Kristensson (2010). Below, we provide a brief overview of the construction and evaluation of such a radial scalar product. We refer the reader to Paper I for a complete pedagogical development.
@@
Note that in Eq. (8), $\xi$ maps $r\in[r_{+},\infty)$ to $\xi\in[0,1)$, thereby compactifying the radial coordinate. In Eq. (9), $\mu(\xi)$ is comprised of factors related to the singular exponents of $\mathcal{L}_{r}$. Its explicit form is given in Eq. (12) of Paper I. In Eq. (10), $\text{C}_{0}$ through $\text{C}_{4}$ are complex valued constants determined by the BH mass $M$, the BH spin $a$, the mode numbers $\{\tilde{\omega},m,l,n\}$, and the field’s spin weight $s$:
@@
Together, Eqs. (8-12) allow the radial eigenvalue relation (Eq. 7) to be rewritten as
@@
${\langle{\mathrm{a}}\,|\,{\mathrm{b}}\rangle}\;=\;\int_{0}^{1}\,\mathrm{a}(\xi)\,\mathrm{b}(\xi)\;\text{W}(\xi)\;d\xi\;,$ (14)
@@
By combining the definition of $\mathcal{L}_{\xi}$ (Eq. 10) with the requirement of self-adjointness (Eq. 15) and then applying integration by parts, it may be shown that
@@
In Eq. (19), $\left<\xi^{j+k}\right$> are called monomial moments Chihara (2011). Using the definition of the weight function, it has been observed (e.g. Cho et al. (2009); Leaver (1986); Abramowitz and Stegun (1964)) that each monomial moment $\left<\xi^{p}\right$> can be evaluated using the Gamma function, $\Gamma(x)$, and the Tricomi confluent hypergeometric function, $U(x,y,z)$,
$\left<\xi^{p}\right>\;=\;e^{\text{B}_{2}}\,\Gamma($ $\text{B}_{0}+p+1)$
$\times\;U(1+\text{B}_{0}+p,-\text{B}_{1},-\text{B}_{2})\;.$ (20)